what are the percentages of the elements C6H12NO2
1
Expert's answer
2020-04-02T08:57:17-0400

Let us consider a mole of "C_6H_{12}NO_2"

mass of one mole of the above compound is "12\times6+12\times1+14+16\times2=130 grams"

Now lets see proportion of each element in it

Carbon- 12*6=72 grams

Hydrogen- 12 grams

Nitrogen- 14 grams

Oxygen- 32 grams

now lets calculate the percentage of each element in mole of the above compound


Carbon= "\frac{72}{130}\times100" =55.38 %


Hydrogen="\frac{12}{130}\times100"= 9.23 %

Nitrogen= "\frac{14}{130}\times100"= 10.77 %


Oxygen= "\frac{32}{130}\times100"= 24.61 %